OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (3, -3) with 0° ≤ θ < 360°

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

r = x^2 + y^2 tan θ = y/x, just make sure you check what quadrant it is (quadrant IV)